# The Fréchet Distance Unleashed: Approximating a Dog with a Frog

### v1

@@ 5 Implementation and experiments
We emphasize however that we have not done low-level optimization of our code for the sake of performance. For instance, better performance would readily follow by implementing our algorithms in C++ or similar low-level languages. Our optimizations tend to be more algorithmic in nature (as described in the rest of paper), by aggressively using simplification, and upper/lower bound computations to guarantee we computed the optimal Fréchet distance.
@@ 5.1 Datasets and hardware
The tests were performed on a Linux system with 64GB memory with Intel i7-11700 CPU, which a decent 16 threads CPU, but far from the fastest hardware currently available.
We use several real-world datasets featured in prior works to demonstrate the effectiveness of our algorithms. Specifically, we use curves taken from the GeoLife dataset zfxml-ggtdu-11 (as used in bkn-wdfpa-21), a stork migration dataset from rktzw-eabgc-18, retrieved from Movebank rktzw-deabg-18 (used in bbknp-klmct-20).
@@ 5.3 Julia
The Julia code is about 5000 lines for the library, and additional 5000 for the examples code. It implements the algorithms described in this paper faithfully, except that there is some additional hacks to handle floating point issues. Julia has, surprisingly easy to use, support for multi-threading (unlike C++ or Python), and we used it here and there in the testing/development.
@@ 5.4.1 Julia vs C++
We compared our implementation performance in Julia to the C++ implementation of Bringmann et al. bkn-wdfpa-21 which seems to be the state of the art. Bringmann et al. also provides code that downloads and generates Fréchet distance tests, which we used ourselves.
We used three data sets. Each data-set is composed of numerous input curves, and then a list of pairs of the curves together with a threshold, and the task at hand is to decider if the Fréchet distance is above/below/equal to the threshold. See Figure 5.2. This task at hand is naturally friendly to simplifications – we implemented a fast simplification data-structure. We iteratively simplify the two curves being compared to progressively finer resolutions, comparing them using the VE-Fréchet distance (monotonizing it). If the comparison is outside the error interval, the result is returned, otherwise, the algorithm goes to the next finer resolution.
Two of the data-sets (characters/Geolife) benefits well from this strategy, and (maybe surprisingly) our Julia code beats the C++ code (which uses completely different algorithm ideas). The third data-set (Sigspatial) seems to be more resistant to this approach, and our code is somewhat slower than C++ code (about 50% slower). One can do massive preprocessing and precompute a simplifications hierarchy for each input curves. With including the preprocessing, the running time more or less match. This is more of a side comment, that storing such precomputed hierarchies might be a good idea if input curves are going to be used repeatedly for performing distance queries.
As for the general C++ vs Julia performance comparison. This seems to be case dependent, with reported cases where either is faster. In our cases, since we did not implement our library in C++, we do not the situation for certain. The significant speed improvement in the case of the Geolife input suggests our algorithm is much faster in some cases where simplification helps. Note, that while our code in Julia is a single thread, it uses automatic garbage collector that runs in a parallel thread.
@@
Input # pairs C++ Julia Julia MT sigspatial $322,000$ 145 263 70 Characters $253,000$ 90 257 67 Geolife $322,000$ $646$ $192$ $69$
Figure 5.2: Julia/C++ comparison. Running times are in seconds. Julia MT stands for the multi-threaded implementation – since multi-threading is very easy in Julia, and the task at hand is easily parallizable, we tested this, but of course this should not be taken as a direct (or remotely fair) comparison to the C++, and is provided for the reader amusement. The tests we performed on a Linux system with 64GB memory with Intel i7-11700 CPU, which a decent 16 threads CPU, but far from the fastest hardware currently available.
@@ 5.4.2 Julia vs Python
Figure 5.4 lists that datasets we sued in comparing the two implementations. See Figure 5.3 and Figure 5.1 for the results (observe that the Julia code was tested for $1.001$ approximation – we have not done the same for the Python code, since it seemed to be significantly slower and numerically unstable in this case). The Python performance on these datasets is competitive with the Julia implementation, although in some cases significantly slower. In general, the Julia implementation is faster (we did not perform enough experiments to decide how much faster), but it is clear the Python implementation is fast and robust enough to be used in practice.
@@ 6 Conclusions and Future Work
In this work, we have demonstrated that our variant of the Fréchet distance is both theoretically efficient and viable in practice. This suggests that our algorithm may be useful in applications where the runtime is dominated by Fréchet distance computations.

### v2

@@ 5 Implementation and experiments
We emphasize however that we have not done low-level optimization of our code for the sake of performance. For instance, better performance might follow by implementing our algorithm in C++ or using squared distances. Our optimizations tend to be more algorithmic in nature (as described in the rest of paper), by aggressively using simplification, and upper/lower bound computations to guarantee we computed the optimal Fréchet distance.
Importantly, because our algorithms use the retractable VE-Fréchet as our building block, the Fréchet morphings our implementation computes look better than what the standard Fréchet matching would give.
@@ 5.1 Datasets and hardware
The tests were all performed on a Linux system with 64GB of memory and an Intel i7-11700 CPU with 16 threads. This system is a fast desktop, but far from the best possible hardware currently available.
We use several real-world datasets featured in prior works to demonstrate the effectiveness of our algorithms. Specifically, we use curves taken from the GeoLife dataset zfxml-ggtdu-11 (as used in bkn-wdfpa-21), a stork migration dataset rktzw-eabgc-18, (retrieved from Movebank rktzw-deabg-18, used in bbknp-klmct-20), and a pigeon flight dataset pgrbvg-pigeon-17. In addition, we use datasets from Bringmann et al. bkn-wdfpa-21 when comparing our implementation to theirs.
@@ 5.3 Julia
The Julia code is about 5000 lines for the library, and additional 5000 for the examples code. It implements the algorithms described in this paper faithfully, with some additional hacks to handle floating point issues. Julia has (surprisingly easy to use) support for multi-threading (unlike C++ or Python), and we used it to speed up various operations.
@@ 5.4.1 Julia vs C++
We compared our implementation performance in Julia to the C++ implementation of Bringmann et al. bkn-wdfpa-21 which seems to be the state of the art. Bringmann et al. also provides code that downloads and generates Fréchet distance tests, which we used ourselves. Of note is that the algorithm of Bringmann et al. uses completely different techniques from our algorithms.
We used three data sets. Each dataset is composed of numerous input curves, and then a list of pairs of the curves together with a threshold, and the task at hand is to decide whether the Fréchet distance is above/below/equal to the threshold. See Figure 5.2. This task at hand is naturally friendly to simplifications – we implemented a fast simplification data-structure. We iteratively simplify the two curves being compared to progressively finer resolutions, comparing them using the (monotonized) VE-Fréchet distance. If the comparison is outside the error interval, the result is returned, otherwise, the algorithm goes to the next finer resolution.
The GeoLife dataset benefits tremendously from the simplification strategy, and (maybe surprisingly) our Julia program is four time faster than the C++ code, see Figure 5.2. The worst dataset for the Julia implementation is the characters dataset where it seems to be about twice slower than the C++ code. This dataset both does not benefit from simplification, and also seem to give rise to many non-monotone morphings, which makes our implementation slower.
The third dataset (Sigspatial) seems to also be somewhat resistant to approximation, and our code is slightly slower than the C++ code. One can do massive preprocessing and precompute a simplifications hierarchy for each input curves. This improves the query process, but makes the preprocessing more expensive. Thus, storing such precomputed hierarchies might be a good idea if input curves are going to be used repeatedly for performing distance queries.
As for the general C++ vs Julia performance comparison. This seems to be case dependent, with reported cases where either is faster. In our cases, since we did not implement our library in C++, we do not know the situation for certain. The significant speed improvement in the case of the GeoLife input suggests our algorithm is much faster in some cases where simplification helps. Note, that while our code in Julia uses a single thread, it uses automatic garbage collector that runs in a parallel thread.
@@
Input # pairs C++ Julia Julia MT sigspatial $322,000$ $145$ $149$ $36$ Characters $253,000$ $90$ $183$ $45$ Geolife $322,000$ $646$ $136$ $63$
Figure 5.2: Julia/C++ comparison. The C++ implementation is from Bringmann et al. bkn-wdfpa-21. Running times are in seconds. Julia MT stands for the multi-threaded implementation – since multi-threading is easy in Julia, and the task at hand is easily parallizable, we tested this, but of course this should not be taken as a direct (or remotely fair) comparison to the C++ implementation, and is provided for the reader amusement. In particular, the multi-threading done is pretty naive, and better performance should be possible by better fine-grained partition of the tasks. The tests we performed on a Linux system with 64GB memory with Intel i7-11700 CPU, with a decent 16 threads CPU, but far from the fastest hardware currently available.
@@ 5.4.2 Julia vs Python
Figure 5.4 lists the datasets we used in comparing the two implementations. See Figure 5.3 and Figure 5.1 for the results (observe that the Julia code was tested for $1.001$ approximation – we have not done the same for the Python code, since it seemed to be significantly slower and numerically unstable in this case). The Python performance on these datasets is competitive with the Julia implementation, although in some cases significantly slower. In general, the Julia implementation is faster (we did not perform enough experiments to decide how much faster), but it is clear that the Python implementation is fast and robust enough to be used in practice.
@@ 6 Conclusions and future work
In this work, we have demonstrated that our variant of the Fréchet distance is both theoretically efficient and viable in practice. We provided standard libraries for the algorithms in Julia and Python. Furthermore, our libraries compute the retractable Fréchet distance/morphing, which seems to perform quite well in practice, and is better at handling noise than the original Fréchet distance. This suggests that our algorithm may be useful in applications where the runtime is dominated by Fréchet distance computations. More generally, we believe the high quality of our implementations, and their easy availability via standard libraries in Julia and Python should make our code and algorithms easily accessible for casual users.